Fama’s Llamas has a weighted average cost of capital of 10.3 percent. The company’s cost of equity is 14 percent, and its pretax cost of debt is 7.5 percent. The tax rate is 21 percent. What is the company’s target debt-equity ratio? (Do not round intermediate calculations and round your answer to 4 decimal places, e.g., .1616.)

please show every step
Answer #1

Weight average cost of capital = Weight of debt * pre tax cost of debt * (1-tax rate) + Weight of equity * cost of equity

=>

weight of equity = 1- weight of debt

=>

let weight of debt is x

=>

10.3% = x * 7.5% * (1-0.21) + (1-x) * 14%

=>

10.3% = x * 5.925% + 14% - 14% x

weight of debt x = 0.45820433436

debt-equity ratio = debt/equity

= 0.45820433436/(1-0.45820433436)

= 0.8457